# The Lax–Wendroff theorem for Patankar-type methods applied to hyperbolic conservation laws

### v1

@@ 3.2 Modified Patankar for difference schemes
The key properties of Patankar-type methods, being unconditionally positivity-preserving and conservative, can be a huge advantage when used in the time integration part of the semidiscretization of a hyperbolic PDE. For instance, they can guarantee to produce nonnegative solutions when approaching zero. This is in contrast with almost all classical time discretization where this bound can either be obtained with constraints on the time-step or it cannot be obtained at all. Negative solutions in such contexts can lead not only to unphysical solutions, but also to NaN in the code and stop the execution. The only major exception is the implicit Euler method, which is unfortunately only first order accurate and very dissipative for hyperbolic conservation laws.
To apply Patankar-type time integration methods to semidiscrete hyperbolic PDEs, it is necessary to find and adequate splitting of the obtained ODE into a PDS. After that, the production and destruction terms are properly weighted. Note that such a splitting is not unique.
@@
Beginning with the PDS, we interpret the fluxes at the cell boundaries as production or destruction terms according to their signs, i.e.,
$p_{i,i+1}(\bm{U}^{n})$ $=d_{i+1,i}(\bm{U}^{n}):=-\min\bigl{\{}g_{i+\frac{1}{2}}^{n},0\bigr{\}},$ (9)
$d_{i,i+1}(\bm{U}^{n})$ $=p_{i+1,i}(\bm{U}^{n}):=\max\bigl{\{}g_{i+\frac{1}{2}}^{n},0\bigr{\}},$
where $\bm{U}^{n}$ denotes the vector containing $U_{i}^{n}$ for $i=1,\dots,N$. We define $p_{i,j},d_{i,j}$ to be equal to zero for $j\neq i+k,\ k\in\{-1,1\}$. Now, we consider a one-stage Patankar-type method interpreted as a non-standard additive Runge–Kutta (NSARK) method with non-standard weights (NSWs) $\delta_{\nu}=\delta_{\nu}(\bm{U}^{n},t_{n},\Delta t)$ [19]. We write it in terms of this PDS in difference form:
@@
$F^{\delta}_{g_{i+\frac{1}{2}}}(\bm{U}^{n}):=d_{i,i+1}(\bm{U}^{n})\delta_{i}-p_{i,i+1}(\bm{U}^{n})\delta_{i+1}=\max\{g^{n}_{i+\frac{1}{2}},0\}\delta_{i}+\min\{g^{n}_{i+\frac{1}{2}},0\}\delta_{i+1},$ (11)
with short notation $F^{n,\delta}_{g_{i+\frac{1}{2}}}=F^{\delta}_{g_{i+\frac{1}{2}}}(\bm{U}^{n}).$ As mentioned earlier, the NSWs $\delta_{\nu}$ depend on the particular method. For example, considering an MPRK scheme as in [21], $\delta_{\nu}$ represent the corresponding Patankar weights, which are $\delta_{i}=\frac{U_{i}^{n+1}}{U_{i}^{n}}$ for the MPE.
@@
It should be noted that the splitting in production and destruction terms is not unique. There are other possibilities that may have advantages regarding, for example, the differentiability of the numerical fluxes. We require that such a splitting results in a conservative PDS consisting of production and destruction terms which are positive on their domain. On the other side, by construction (9), there always exists a splitting.
[…]
@@
for $k=1,\dots,s$, where new NSWs $\gamma_{\nu}^{(k)}=\gamma_{\nu}^{(k)}(\bm{U}^{n},t_{n},\Delta t)$ are introduced for the stages $U_{i}^{(k)}$. We define the new numerical flux based on $F^{\delta}_{g_{i+\frac{1}{2}}}(\bm{U}^{n})$ from (11) as
$\bar{F}^{n,\delta}_{g_{i+\frac{1}{2}}}:=\sum\nolimits_{r=1}^{s}b_{r}F^{\delta}_{g_{i+\frac{1}{2}}}(\bm{U}^{(r)}).$ (13)
@@
This approach can also be extended to systems of conservation laws. There, we apply the modified Patankar-trick on every component of the vector $\mathbf{u}$ of the initial value problem that requires positivity, and we consider t he numerical flux for every such component as a scalar function $g$. With that it becomes evident that this approach can be applied also for different space-discretizations as long as they can be written in flux formulation [1].
@@
The Modified Patankar RK can be designed also when some of the $\mathbf{A}$ and $\mathbf{b}$ coefficients are negative, see [21], by simply changing the role of production and destruction for that particular term in that stage. This has been applied systematically to design arbitrary high order Modified Patankar methods as the MPDeC [33] that will be used in the numerical section.
@@
The difference schemes framework (2) is very general and can encompass various spatial discretization. This includes also various high order discretizations as ENO, WENO finite volume [10, 9] and finite difference [17] reconstructions schemes. A slightly different formulation should be introduced to include also other high order discretizations as continuous and discontinuous Galerkin Finite Element methods [32, 44], where the PDS is either written only on the interfaces of the cells or there is more freedom in the choice of the PD terms between the degrees of freedom inside each cell.
4 Lax–Wendroff theorem for MP-difference schemes
The aim of this section is to adapt the proof of Theorem 1 to Patankar-type schemes in the form (12). To achieve this, we need to control the nonlinear weights, resulting in the extra assumption, that the numerical solution has a bounded total variation also in time (classical Lax–Wendroff requires only bounded total variation in space). So, let us introduce the concept of total time variation (TTV).
@@ 4 Lax–Wendroff theorem for MP-difference schemes
The TTV over a domain $[a,b]\times[0,T]$ of a function $U\in L^{1}_{\text{loc}}(\mathbb{R}\times\mathbb{R}^{+})$ is
@@
On a domain $[a,b]\times[0,T]$ for a discrete function $\{U_{i}^{n}\}_{i,n}=\{U(x_{i},t_{n})\}_{i,n}$ for $i=1,\dots,N$ and $n=0,\dots,n_{T}$ with $t^{0}=0$ and $t^{n_{T}}=T$, the TTV can be equivalently defined using a piece-wise constant reconstruction of the discrete function on the space-time grid, obtaining the formula
@@
[…]
Proof 1.
@@
$\delta^{\{l\}}(x,t):=\frac{U^{\{l\}}(x,t+\Delta t_{\{l\}})}{U^{\{l\}}(x,t)}=1+\frac{U^{\{l\}}(x,t+\Delta t_{\{l\}})-U^{\{l\}}(x,t)}{U^{\{l\}}(x,t)}.$
Now, we can say that $\lim_{l\to\infty}U^{\{l\}}(x,t+\Delta t_{\{l\}})-U^{\{l\}}(x,t)=0$ a. e. because the numerical solutions are TTV bounded uniformly for all $l$ and $x$ and the solutions are converging to a weak solution $u$ in $L^{1}_{\text{loc}}$.
The arguments of this proof are the same used in space for the Lax–Wendroff theorem in the proof of [28].
@@
Consider a sequence of step sizes $\Delta x_{\{l\}}>0$ with $\lim_{l{\rightarrow\infty}}\Delta x_{\{l\}}=0$. Let $\left(U^{\{l\}}\right)_{l}$ be a sequence of piece-wise constant functions constructed of numerical solutions obtained by a MPE difference scheme (10) consistent with (1) with time step size $\Delta t_{\{l\}}=\lambda\Delta x_{\{l\}},\lambda\equiv\mathrm{const}.$ If for any domain $[a,b]\times[0,T]$ the total variation $TV\left(U^{\{l\}}([a,b],t)\right)$ from (15) is uniformly bounded for all $l$ and $t\in[0,T]$, the total time variation $TTV\left(U^{\{l\}}(x,[0,T])\right)$ is uniformly bounded for all $l$ and $x\in[a,b]$ and if $U^{\{l\}}\xrightarrow{l\rightarrow\infty}u$ in $L^{1}_{\mathrm{loc}}(\mathbb{R}\times[0,T))$, then $u$ is a weak solution of the conservation law (1).
Proof 2.
@@
for all $\varphi\in C_{0}^{1}([a,b]\times[0,T])$. In what follows, we omit to write the dependency on the refinement represented by the index $l$ if it is clear from the context and extend the numerical solution, if needed, periodically to the domain (,)×$(-\infty,\infty)\times[0,\infty)$.
Multiplying the numerical method (10) by any test function $\varphi\in C_{0}^{1}$ and summation over all $i$ and $n\geq 0$ yields
$\sum_{n=0}^{\infty}\sum_{i=-\infty}^{\infty}\varphi(x_{i},t_{n})(U_{i}^{n+1}-U_{i}^{n})=\frac{\Delta t}{\Delta x}\sum_{n=0}^{\infty}\sum_{i=-\infty}^{\infty}\varphi(x_{i},t_{n})\left[F^{\delta}_{g_{i+\frac{1}{2}}}(\bm{U}^{n})-F^{\delta}_{g_{i-\frac{1}{2}}}(\bm{U}^{n})\right],$
whereas using summation by parts, the fact $\varphi$ has a compact support, multiplying by $\Delta x$ and rearranging leads to
@@
Now we show that this equation coincides with the desired assertion (16) as we pass to the limit $\lim_{l{\rightarrow\infty}}\Delta x_{\{l\}}=0$ with $\Delta t_{\{l\}}=\lambda\Delta x_{\{l\}},\lambda\equiv\mathrm{const}$. For the first and the last term of the previous equation, apply the classical proof of Lax and Wendroff, as done in [28], i.e., as $l\rightarrow\infty$, the $L^{1}_{\mathrm{loc}}$ convergence of $U^{\{l\}}$ to $u$, together with the smoothness of $\varphi$, ensures the convergence of the first summand to $\int_{0}^{\infty}\int_{-\infty}^{\infty}\varphi_{t}(x,t)u(x,t)\operatorname{d}\!x$, whereas the choice of the initial data $U_{i}^{0}=\psi_{i}$ enables us to conclude the convergence of the RHS to $-\int_{-\infty}^{\infty}\varphi(x,0)u(x,0)\operatorname{d}\!x$. What is left to prove is the limit for the middle term, in particular, we are interested showing the convergence of the new numerical flux to the physical flux and therefore in estimating
@@
for $l\to\infty$ where $F^{\delta,\{l\}}_{g}((\mathbf{x}_{i},t_{n})$ is a piece-wise constant reconstruction of the values $F^{\delta}_{g_{i+\frac{1}{2}}}(\bm{U}^{n})$ from (11) by using
@@
for $i=1,\dots,N_{l}$ and $n=0,\dots,n_{T,\{l\}}$ with $\mathbf{x}_{i}=(x_{i}-p\Delta x,\dots,x_{i}+q\Delta x)$. Adding and subtracting
@@
where the first two addends vanish a.e. for $l\rightarrow\infty$ due to Lemma 2.
Finally, as only the original numerical flux $g$ appears in the remaining term $\left\lvert g(U^{\{l\}}(\mathbf{x}_{i},t_{n}))-f(\bm{U}^{\{l\}}(x_{i},t_{n}))\right\rvert$, we can follow the reasoning for Lax–Wendroff [28]. Consistency of $g$ is provided, i.e., we have the estimate
@@
with a Lipschitz constant $K$, where the total variation boundness (TVB) property of $U^{\{l\}}(\cdot,t)$ leads to
@@
as $l\rightarrow\infty$ for almost every value of $x$. Altogether we proved the a. e. convergence of the numerical flux function $F^{\delta,\{l\}}_{g}(x_{i},t_{n})$ to the physical flux function $f(U^{\{l\}}(x_{i},t_{n}))$. This gives, together with some standard estimates, the convergence of (17) to the weak form (16) and shows, since it is true for any test function $\varphi\in C_{0}^{1}$, that $u$ is indeed a weak solution.

### v2

@@ 3.2 Modified Patankar for difference schemes
The key properties of Patankar-type methods, being unconditionally positivity-preserving and conservative, can be a huge advantage when used as the time integration scheme of the semidiscretization of a hyperbolic PDE. For instance, they can guarantee to produce nonnegative solutions when approaching zero. This is in contrast with almost all classical time discretizations where this bound can either be obtained with constraints on the time-step or it cannot be obtained at all. Negative solutions in such contexts can lead not only to unphysical solutions, but also to NaN in the code and stop the execution. The only major exception is the implicit Euler method, which is unfortunately only first order accurate and very dissipative for hyperbolic conservation laws.
To apply Patankar-type time integration methods to semidiscrete hyperbolic PDEs, it is necessary to find and adequate splitting of the obtained ODE into a PDS. After that, the production and destruction terms are properly weighted. Note that such a splitting is not unique, see Remark 2.
@@
Starting with the PDS, we interpret the fluxes at the cell boundaries as production or destruction terms according to their signs, i.e.,
$p_{i,i+1}(\bm{U}^{n})$ $=d_{i+1,i}(\bm{U}^{n})\mathrel{\mathop{\ordinarycolon}}=-\min\bigl\{g_{i+\frac{1}{2}}^{n},0\bigr\},$ (9)
$d_{i,i+1}(\bm{U}^{n})$ $=p_{i+1,i}(\bm{U}^{n})\mathrel{\mathop{\ordinarycolon}}=\max\bigl\{g_{i+\frac{1}{2}}^{n},0\bigr\},$
where $\bm{U}^{n}$ denotes the vector containing $U_{i}^{n}$ for $i=1,\dots,N$. With that, we can set $p_{i,j},d_{i,j}$ to zero for $j\notin\{i-1,i+1\}\subseteq\{1,\dotsc,N\}$. Now, we consider a one-stage Patankar-type method interpreted as a non-standard additive Runge–Kutta (NSARK) method with non-standard weights (NSWs) $\delta_{\nu}=\delta_{\nu}(\hat{\bm{U}}^{n},t_{n},\Delta t)$ [20]. We write it in terms of this PDS in difference form:
@@
$F^{\delta}_{g_{i+\frac{1}{2}}}(\bm{U}^{n})\mathrel{\mathop{\ordinarycolon}}=d_{i,i+1}(\bm{U}^{n})\delta_{i}-p_{i,i+1}(\bm{U}^{n})\delta_{i+1}=\max\{g^{n}_{i+\frac{1}{2}},0\}\delta_{i}+\min\{g^{n}_{i+\frac{1}{2}},0\}\delta_{i+1}$ (11)
and the short notation $F^{n,\delta}_{g_{i+\frac{1}{2}}}=F^{\delta}_{g_{i+\frac{1}{2}}}(\bm{U}^{n}).$ As mentioned earlier, the NSWs $\delta_{\nu}$ depend on the particular method. For example, considering an MPRK scheme as in [23], $\delta_{\nu}$ represent the corresponding Patankar weights, which are $\delta_{i}=\frac{U_{i}^{n+1}}{U_{i}^{n}}$ for the MPE.
@@
First of all, with the construction (9), there always exists a splitting of the conservation law into a conservative PDS. However, it should be noted that the splitting in production and destruction terms is not unique as, in general, one could add a positive constant to all production and destructions $p^{\text{new}}_{ij}\coloneqq p^{\text{old}}_{ij}+c$ for all $i\neq j$, and $d^{\text{new}}_{ij}\coloneqq p^{\text{new}}_{ji}$, leading to another valid splitting of the original PDS for any $c\geq 0$.
Furthermore, there might be even multiple non-trivial splittings, e. g., for the inviscid Burgers’ equation using an upwind FV discretization, where $U_{i}\geq 0,\,f(U_{i})=\frac{1}{2}U_{i}^{2}\geq 0$, we obtain
$\frac{\mathrm{d}U_{i}}{\mathrm{dt}}=$ $-\frac{1}{\Delta x}\left(f(U_{i})-f(U_{i-1})\right)=\frac{1}{\Delta x}f(U_{i-1})-\frac{1}{\Delta x}f(U_{i})$
$=$ $\hphantom{+}\frac{1}{\Delta x}\left(-(f(U_{i})-f(U_{i-1}))^{-}+f(U_{i})\right)\;-\frac{1}{\Delta x}(f(U_{i})-f(U_{i-1}))^{+}\;$
$+\frac{1}{\Delta x}\left((f(U_{i+1})-f(U_{i}))^{-}-f(U_{i+1})\right)+(f(U_{i+1})-f(U_{i}))^{+},$
where we introduced $\tau^{+}\coloneqq\max(\tau,0)$, $\tau^{-}\coloneqq\min(\tau,0)$ for a scalar function $\tau$, i. e., $\tau^{+}+\tau^{-}=\tau$. Here, both re-formulations may be interpreted as different conservative PDS splittings satisfying $p_{i,j}=0$ for $j\notin\{i-1,i+1\}\subseteq\{1,\dotsc,N\}$ , $d_{i,j}=p_{j,i}$ for all $i,j$, and
$p_{i,i-1}(\bm{U})=\frac{1}{\Delta x}f(U_{i-1}),\qquad d_{i,i+1}=-\frac{1}{\Delta x}f(U_{i}),\qquad p_{i,i+1}(\bm{U})=d_{i,i-1}=0\qquad\text{ or }$
$p_{i,i-1}(\bm{U})$ $=\frac{1}{\Delta x}\left(-(f(U_{i})-f(U_{i-1}))^{-}+f(U_{i})\right),\qquad d_{i,i-1}(\bm{U})=\frac{1}{\Delta x}(f(U_{i})-f(U_{i-1}))^{+},$
$d_{i,i+1}(\bm{U})$ $=-\frac{1}{\Delta x}\left((f(U_{i+1})-f(U_{i}))^{-}-f(U_{i+1})\right),\qquad p_{i,i+1}(\bm{U})=\frac{1}{\Delta x}(f(U_{i+1})-f(U_{i}))^{+},$
respectively. It is worth noting that the first splitting corresponds to (9) and possesses differentiable production and destruction terms in this example, while the second splitting does not. Also, by comparing $p_{i,i+1}$ it becomes evident that the two splittings differ. In what follows, we require that any splitting of the PDE results in a conservative PDS consisting of production and destruction terms which are non-negative on their domain.
Although, we first introduced one stage methods, also an $s$-stage Patankar-type difference scheme based on an explicit RK scheme with $\bm{A}=(a_{kr})_{k,r=1,\dots,s}\geq\b${0} and weights $b_{1},\dots,b_{s}\geq 0$ can be devised (using $p_{i,j}=d_{j,i}$) by
@@
for $k=1,\dots,s$, where new NSWs $\gamma_{\nu}^{(k)}=\gamma_{\nu}^{(k)}(\hat{\bm{U}}^{n},t_{n},\Delta t)$ are introduced for the stages $U_{i}^{(k)}$. We define the new numerical flux based on $F^{\delta}_{g_{i+\frac{1}{2}}}(\bm{U}^{n})$ from (11) as
$\bar{F}^{n,\delta}_{g_{i+\frac{1}{2}}}\mathrel{\mathop{\ordinarycolon}}=\sum\nolimits_{r=1}^{s}b_{r}F^{\delta}_{g_{i+\frac{1}{2}}}(\bm{U}^{(r)}).$ (13)
@@
This approach can also be extended to systems of conservation laws. There, we apply the MP trick on every component of the vector $\mathbf{u}$ of the initial value problem that requires positivity, and we consider the numerical flux for every such component as a scalar function $g$. With that it becomes evident that this approach can be applied also for different space-discretizations as long as they can be written in flux formulation [1].
@@
The MPRK can be designed also when some of the $\mathbf{A}$ and $\mathbf{b}$ coefficients are negative, see [23], by simply changing the role of production and destruction for that particular term in that stage. This has been applied systematically to design arbitrary high order MP methods as the MPDeC [35] that will be used in the numerical section.
@@
The difference schemes framework (2) is very general and can encompass various spatial discretizations. This includes also various high order discretizations as Essentially Non-Oscillatory (ENO), WENO, FV [12, 11] and finite difference [18] reconstructions schemes. A slightly different formulation should be introduced to include also other high order discretizations as continuous and DG FE methods [34, 46], where the PDS is either written only on the interfaces of the cells or there is more freedom in the choice of the PD terms between the degrees of freedom inside each cell.
4 Lax–Wendroff theorem for MP difference schemes
The aim of this section is to adapt the proof of Theorem 1 to Patankar-type schemes in the form (12). To achieve this, we need to control the nonlinear weights, resulting in the extra assumption that the numerical solution has a bounded total variation also in time (classical Lax–Wendroff requires only bounded total variation in space). So, let us introduce the concept of total time variation (TTV).
@@ 4 Lax–Wendroff theorem for MP difference schemes
The TTV over a domain $a,b]\times[0,T]$ of a function $U\in L^{1}_{\text{loc}}(\mathbb{R}\times\mathbb{R}^{+})$ is
@@
On a domain $a,b]\times[0,T]$ for a discrete function $\{U_{i}^{n}\}_{i,n}=\{U(x_{i},t_{n})\}_{i,n}$ for $i=1,\dots,N$ and $n=0,\dots,n_{T}$ with $t_{0}=0$ and $t_{n_{T}}=T$, the TTV can be equivalently defined using a piece-wise constant reconstruction of the discrete function on the space-time grid, obtaining the formula
@@
Remark 7 (TTV uniform boundedness).
The TTV uniform boundedness, i.e., having a bound for the TTV of an approximate solution while the mesh is refined, will be used as a requirement in the next sections in many theorems. As for the TV boundedness in space, it is related to the non-oscillatory behavior of the numerical solution in time and, hence, their stability. Moreover, we believe that there is a link between the TV boundedness and the TTV boundedness for hyperbolic problems, though it is not strainghtforward to prove it. During numerical results, we have seen that TTV boundedness seems naturally achieved (with $\text{TTV}(U^{num}(x))$ converging to the exact $\text{TTV}(U^{ex}(x))$) when a stable CFL number is used. On the contrary, when oscillations arise in the space discretization, they also propagate in time, leading to not converging $\text{TTV}(U^{num}(x))$.
@@
Let $U^{\{l\}}(x_{i},t_{n})$ be the approximation of problem (1) at the space $x^{\{l\}}_{i}$ and time $t_{n}^{\{l\}}$ for a mesh discretization $\Delta x_{\{l\}}\sim\Delta t_{\{l\}}\to 0$ for $l\to\infty$. For every $l$, consider $U^{\{l\}}(x,t)$ as a piece-wise constant reconstruction of the numerical solution $U^{\{l\}}(x_{i},t_{n})$. Let $\text{TTV}(U^{\{l\}}(x,[0,T]))$ be uniformly bounded for all $x\in[a,b]$ and for all $l\in\mathbb{N}$. Suppose that $U^{\{l\}}\to u$ in $L^{1}_{\text{loc}}([a,b]\times[0,T])$ with $u$ positive. Defining $\delta^{\{l\}}(x,t)$ as the piece-wise constant reconstruction of the MPE weights $\delta^{\{l\}}(x_{i},t_{n})=U^{\{l\}}(x_{i},t_{n+1})/U^{\{l\}}(x_{i},t_{n})$ for all $i$ and $n$. Then, $\delta^{\{l\}}(x,t)\to 1$ almost everywhere (a.e.) in $a,b]\times[0,T]$.
Proof.
@@
$\delta^{\{l\}}(x,t)\mathrel{\mathop{\ordinarycolon}}=\frac{U^{\{l\}}(x,t+\Delta t_{\{l\}})}{U^{\{l\}}(x,t)}=1+\frac{U^{\{l\}}(x,t+\Delta t_{\{l\}})-U^{\{l\}}(x,t)}{U^{\{l\}}(x,t)}.$
Now, we can say that $\lim_{l\to\infty}U^{\{l\}}(x,t+\Delta t_{\{l\}})-U^{\{l\}}(x,t)=0$ a. e. because the numerical solutions are TTV bounded uniformly for all $l$ and $x$ and the denominator, $U^{\{l\}}$ is converging to a weak solution $u$ in $L^{1}_{\text{loc}}$. ∎
The arguments of this proof are the same used in space for the Lax–Wendroff theorem in the proof of [30]. These follow from some classical analysis results using the compactness of bounded total variation functions in $L^{1}_{loc}$, the fact that in 1D functions of bounded variation have at most countably many discontinuities and the Helly’s selection theorem to deduce that bounded sequences of 1D bounded total variation functions are precompact for pointwise convergence almost everywhere [2].
@@
Consider a sequence of step sizes $\Delta x_{\{l\}}>0$ with $\lim_{l{\rightarrow\infty}}\Delta x_{\{l\}}=0$. Let $\left(U^{\{l\}}\right)_{l}$ be a sequence of piece-wise constant functions constructed of numerical solutions obtained by a MPE difference scheme (10) consistent with (1) with time step size $\Delta t_{\{l\}}=\lambda\Delta x_{\{l\}},\lambda\equiv\mathrm{const}.$ If for any domain $a,b]\times[0,T]$ the total variation $TV\left(U^{\{l\}}([a,b],t)\right)$ from (15) is uniformly bounded for all $l$ and $t\in[0,T]$, the total time variation $TTV\left(U^{\{l\}}(x,[0,T])\right)$ is uniformly bounded for all $l$ and $x\in[a,b]$ and if $U^{\{l\}}\xrightarrow{l\rightarrow\infty}u$ in $L^{1}_{\mathrm{loc}}(\mathbb{R}\times[0,T))$, then $u$ is a weak solution of the conservation law (1).
Proof.
@@
for all $\varphi\in C_{0}^{1}([a,b]\times[0,T])$. In what follows, we omit to write the dependency on the refinement represented by the index $l$ if it is clear from the context.
Multiplying the numerical method (10) by any test function $\varphi\in C_{0}^{1}$ and summing over all $i$ and $n\geq 0$ yields
$\sum_{n=0}^{\infty}\sum_{i=-\infty}^{\infty}\varphi(x_{i},t_{n})(U_{i}^{n+1}-U_{i}^{n})=\frac{\Delta t}{\Delta x}\sum_{n=0}^{\infty}\sum_{i=-\infty}^{\infty}\varphi(x_{i},t_{n})\left[F^{\delta}_{g_{i+\frac{1}{2}}}(\bm{U}^{n})-F^{\delta}_{g_{i-\frac{1}{2}}}(\bm{U}^{n})\right].$
We note that the value of, e. g., $U_{i}^{n}$ does not have an effect for $(x_{i},t_{n})\notin\operatorname{supp}(\varphi)\subseteq[a,b]\times[0,T]$. Using summation by parts, the fact $\varphi$ has a compact support, multiplying by $\Delta x$ and rearranging leads to
@@
Now we show that this equation coincides with the desired assertion (16) as we pass to the limit $\lim_{l{\rightarrow\infty}}\Delta x_{\{l\}}=0$ with $\Delta t_{\{l\}}=\lambda\Delta x_{\{l\}},\lambda\equiv\mathrm{const}$. For the first and the last term of the previous equation, the classical proof of Lax and Wendroff should be applied, as outlined in [30]. This means that as $l\rightarrow\infty$, the assumed $L^{1}_{\mathrm{loc}}$ convergence of $U^{\{l\}}$ to $u$, together with the smoothness of $\varphi$, ensures the convergence of the first summand to $\int_{0}^{\infty}\int_{-\infty}^{\infty}\varphi_{t}(x,t)u(x,t)\operatorname{d}\!x$. In addition, the choice of the initial data $U_{i}^{0}=\psi_{i}$ enables us to conclude the convergence of the RHS to $\int_{-\infty}^{\infty}\varphi(x,0)u(x,0)\operatorname{d}\!x$. The remaining task is to ascertain the limit for the middle term. In particular, we are interested in showing the convergence of the new numerical flux to the physical flux and therefore in estimating
@@
for $l\to\infty$ where $\mathbf{x}_{i}=(x_{i}-p\Delta x,\dots,x_{i}+q\Delta x)$. Here, $F^{\delta,\{l\}}_{g}(\mathbf{x}_{i},t_{n})$ is a piece-wise constant reconstruction of the values $F^{\delta}_{g_{i+\frac{1}{2}}}(\bm{U}^{n})$ given in (11) by substituting
@@
for $i=1,\dots,N_{l}$ and $n=0,\dots,n_{T,\{l\}}$ for $g_{i+\frac{1}{2}}^{n}$. Adding and subtracting
@@
where the first two addends vanish a.e. as $l\rightarrow\infty$ due to the convergence of the MPE weights based on the TTV boundedness property, as stated in Lemma 2.
Finally, as only the original numerical flux $g$ appears in the remaining term $\left\lvert g(U^{\{l\}}(\mathbf{x}_{i},t_{n}))-f(\bm{U}^{\{l\}}(x_{i},t_{n}))\right\rvert$, we can follow the reasoning of the Lax–Wendroff theorem of [30]. Consistency of $g$ is provided, i.e., we have the estimate
@@
with a Lipschitz constant $K$, where the total variation boundedness (TVB) property of $U^{\{l\}}(\cdot,t)$ leads to
@@
as $l\rightarrow\infty$ for almost every value of $x$. Altogether we proved the a. e. convergence of the numerical flux function $F^{\delta,\{l\}}_{g}(\mathbf{x}_{i},t_{n})$ to the physical flux function $f(U^{\{l\}}(x_{i},t_{n}))$. This gives, together with some standard estimates, the convergence of (17) to the weak form (16) and shows, since it is true for any test function $\varphi\in C_{0}^{1}$, that $u$ is indeed a weak solution.